The cost of a membership to a health club has an initiation fee and monthly dues. If the club charges $100 for 2 months and $200 for 6 months, then find the initiation cost and the monthly dues fees.
Answer by robertb(5830)

Let i = initiation fee, m = monthly fee.
Then we get the system
i + 2m = 100 <--(1)
i + 6m = 200 <--(2)
Subtracting the top equation from the bottom equation, we get
4m = 100, or m = $25, the monthly fee.
The initiation fee is i + 2*25 = 100,
==> i = 100 - 50 = $50